OpenStudy (anonymous):

f(1-x)-x*f(1/x)=x-2 find ''x'' ?

OpenStudy (anonymous):

anyidea what is f(x)?

OpenStudy (anonymous):

no.just find f(x)

OpenStudy (anonymous):

f(x) means function of x or read f function of x

OpenStudy (anonymous):

did you copy the prob correctly?

OpenStudy (anonymous):

i think we can say f(1-x)=(xf(1/x))+(x-2)

OpenStudy (anonymous):

1-x=t ==> x=1-t f(t)=(1-t)(f(1/(1-t))-(t+1) but i stuck in here

OpenStudy (anonymous):

re u sure u can do that in function? "1-x=t ==> x=1-t"

OpenStudy (anonymous):

we learn this in school let me show an example : f(3x-2)=5x, what is f(x)? answer: first 3x-2=t then x=(t+2)/3 at end : f(t)=5*(t+2)/3 then change ''t'' to x f(x)=5*(x+2)/3

OpenStudy (anonymous):

cool, learnt sumthing new